What is the moment of inertial of a 3kg uniform rod with length 2 meters and radius


20cm, if the axis of rotation is located at (a) one end of the rod, (b) at the center of the


rod and (c) through the center?



1
Expert's answer
2022-04-17T17:08:33-0400

See the formulas for the moment of inertia here: https://en.wikipedia.org/wiki/List_of_moments_of_inertia.


Let "L=2m" be the length of the rod, "m=3kg" be its mass, and "r = 20cm = 0.2m" be its radius. Then the moment of inertial will be the following.


a) "I = \dfrac13 mL^2 = \dfrac{3kg\cdot (2m)^2}{3} = 4kg\cdot m^2"

b) "I = \dfrac{1}{12} mL^2 = \dfrac{3kg\cdot (2m)^2}{12} = 1kg\cdot m^2"

c) "I = \dfrac{1}{2} mr^2 = \dfrac{3kg\cdot (0.2m)^2}{2} = 0.06kg\cdot m^2"